Find the laplace transform of t2?

1
Expert's answer
2021-07-19T05:32:13-0400

Taking into account that the laplace transform of "t^n" is "L(t^n)=\frac{n!}{s^{n+1}}", we conclude that the laplace transform of "t^2" is "L(t^2)=\frac{2}{s^{3}}."